# WARP-LCA: Efficient Convolutional Sparse Coding with Locally Competitive Algorithm

### v1

@@ Abstract
The locally competitive algorithm (LCA) can solve sparse coding problems across a wide range of use cases. Recently, convolution-based LCA approaches have been shown to be highly effective for enhancing robustness for image recognition tasks in vision pipelines. To additionally maximize representational sparsity, LCA with hard-thresholding can be applied. While this combination often yields very good solutions satisfying an $\ell_{0}$ sparsity criterion, it comes with significant drawbacks for practical application: (i) LCA is very inefficient, typically requiring hundreds of optimization cycles for convergence; (ii) the use of a hard-thresholding results in a non-convex loss function, which might lead to suboptimal minima. To address these issues, we propose the Locally Competitive Algorithm with State Warm-up via Predictive Priming (WARP-LCA), which leverages a predictor network to provide a suitable initial guess of the LCA state based on the current input. Our approach significantly improves both convergence speed and the quality of solutions, while maintaining and even enhancing the overall strengths of LCA. We demonstrate that WARP-LCA converges faster by orders of magnitude and reaches better minima compared to conventional LCA. Moreover, the learned representations are more sparse and exhibit superior properties in terms of reconstruction and denoising quality as well as robustness when applied in deep recognition pipelines. Furthermore, we apply WARP-LCA to image denoising tasks, showcasing its robustness and practical effectiveness. Our findings confirm that the naive use of LCA with hard-thresholding results in suboptimal minima, whereas initializing LCA with a predictive guess results in better outcomes. This research advances the field of biologically inspired deep learning by providing a novel approach to convolutional sparse coding.
@@ 1 Introduction
Sparse coding has deep roots in neuroscience. The idea that sensory systems learn and represent the statistics of natural scenes date back to Barlow’s efficient coding hypothesis (Barlow et al., 1961). Barlow reasoned that redundancy reduction is key, since information retrieval in the sensory stream is akin to finding the needle in the haystack. Sparse coding then originated from the insight that sparse, redundant representations may be useful to make the statistical components of sensory information explicit (Field, 1994; Barlow, 2001). Within this context, the sparse coding algorithm, introduced by (Olshausen and Field, 1996), has emerged as a paradigm to successfully model and predict response patterns of the primary visual cortex (Lee et al., 2006, 2007; Beyeler et al., 2019; Ecke et al., 2021).
@@
The human cortex is a very homogeneous structure, and it has been hypothesized that it performs the same operation everywhere (Barlow, 1987). The claim that sparse coding, at least to some degree, models general cortical function, reflects in a manifold of applications in which the algorithm has proven useful. It has early on been shown that the learned features are useful for inference (Rigamonti et al., 2011), especially when unsupervised learning can be leveraged, like in stereo vision (Lundquist et al., 2016, 2017; Ecke et al., 2021) or in scenarios with poorly labeled data (Zhang et al., 2017). The capabilities for unsupervised learning are most prominent in multi layer sparse coding networks, which are capable of learning abstract concepts (Kim et al., 2018; Zhang et al., 2019; Dibbo et al., 2023). In addition, sparse coding has proven at least state-of-the-art performance in removing noise and other corruptions from images (Lecouat et al., 2020). More recently, Teti et al. (2022) has shown that applying sparse coding as a pre-processing stage in image recognition pipelines yields state-of-the-art robustness to common corruptions and adversarial attacks.
@@
A significant challenge for applications with sparse coding is the computational expense of the inference algorithm. Finding sparse coefficients for each input sample requires an expensive iterative optimization process via gradient descent. One approach to resolve poor execution speed is the implementation on neuromorphic hardware. Rozell et al. (2008) introduced the Locally Competitive Algorithm (LCA) as a biologically plausible model for sparse coding. Indeed, the implementation of the LCA on Intel Loihi has proven the strongest gain in energy efficiency and execution speed against the implementation on conventional hardware, as compared to any other implemented algorithm (Davies et al., 2021; Henke et al., 2022; Parpart et al., 2023). Considerable research has been directed towards developing efficient optimization algorithms for sparse coding on conventional hardware. Notable contributions include the works of Li and Osher (2009); Mairal et al. (2009); Beck and Teboulle (2009), and Vonesch and Unser (2007). Gregor and LeCun (2010) contributed significantly to the field by introducing the Learned ISTA (LISTA) algorithm, which has spawned a family of model-based sparse encoders (notable models include AMP-Net Zhang et al. (2020), FISTA-Net Xiang et al. (2021) and ISTA-Net Zhang and Ghanem (2018).
@@
Sparse coding is typically associated with $\ell_{1}$ norm regularization due to its convexity, which facilitates a tractable optimization process (Zhang et al., 2015). However, for achieving maximum sparsity, $\ell_{0}$-like regularization is preferred (Paiton, 2019). Despite its non-convex nature, which introduces computational challenges, $\ell_{0}$ regularization directly targets sparsity in the solution (Nguyen et al., 2019). Accelerated methods such as Learned Iterative Hard Thresholding (L-IHT) and Hard Thresholding Pursuit (HTP) have been developed to address $\ell_{0}$-like convergence problems. Nonetheless, the Locally Competitive Algorithm (LCA) presents a distinctive case. While the discrete version of LCA implementing $\ell_{1}$ sparsity aligns mathematically with ISTA  (Balavoine et al., 2015), employing an $\ell_{0}$-like cost function in LCA diverges significantly in outcomes when compared to greedy methods such as Basis Pursuit (BP) (Rozell et al., 2007, 2008).
@@
Our main contribution of this paper is the development and validation of the Locally Competitive Algorithm with State Warm-up via Predictive Priming (WARP-LCA). This novel approach leverages a predictor network to provide an initial guess of the LCA state based on the current input, significantly improving convergence speed and solution quality. Specifically, WARP-LCA addresses the inefficiencies and suboptimal solutions associated with LCA using hard-thresholding ($\ell_{0}$). By incorporating predictive priming, WARP-LCA dramatically reduces the number of optimization cycles required for convergence. Furthermore, WARP-LCA achieves superior minima, yielding more sparse and robust representations that enhance reconstruction and denoising quality, particularly in deep recognition pipelines. Our experiments, in the domain of image denoising and comparing WARP-LCA directly to LCA, confirm that WARP-LCA not only overcomes the limitations of traditional hard-thresholding LCA but also advances the field of biologically inspired deep learning by providing a novel and efficient method for convolutional sparse coding.
@@ 2.1 Convolutional Sparse Coding
The inference process involves estimating the sparse coefficients $a$ for a given data point $x$. This is typically achieved by minimizing an energy function that combines a reconstruction error term with a sparsity-inducing regularization term. A common formulation uses $\ell_{p}$ regularization to enforce sparsity (Tibshirani, 1996).
@@
Learning in convolutional sparse coding involves updating the convolutional filters $\Phi$ based on the inferred sparse coefficients across a batch of data points. This is typically done through gradient descent on the energy function with respect to $\Phi$. Although the joint optimization of $\Phi$ and $a$ is non-convex, iterative approaches often converge to useful solutions, especially in natural signal domains. In addition to convolutional sparse coding, other widely known approaches for learning dictionaries include Principal Component Analysis (PCA) and Independent Component Analysis (ICA). PCA learns an orthogonal dictionary that captures the directions of maximum variance in the data. ICA, on the other hand, seeks to represent data as linear combinations of statistically independent components, which is useful for blind source separation and feature extraction (Hyvärinen and Oja, 2000).
@@ 2.2 Locally Competitive Algorithm (LCA)
The Locally Competitive Algorithm (LCA), introduced by Rozell et al. Rozell et al. (2008), represents a biologically plausible approach to sparse coding, where the dynamics of the neuronal population can be conceptualized as a system of coupled differential equations that govern the temporal evolution of neuronal activities. Unlike iterative thresholding methods or matching pursuit, the LCA exhibits a ’charging circuit behavior’ where neurons dynamically compete to represent the input.
@@
Rozell et al. (2008) introduced a generalized thresholding function that enables the Locally Competitive Algorithm (LCA) to tackle various sparse coding problems by adjusting its parameters. This function can interpolate between $\ell_{1}$ and $\ell_{0}$ regularization, providing a spectrum of sparsity constraints. The generalized thresholding function is expressed as:
@@
Specifically, for $\ell_{0}$ regularization—which aims to minimize the number of non-zero coefficients—the thresholding function approaches a hard threshold as $\gamma$ becomes large. This effectively turns the generalized thresholding function into:
@@
The initial values of the states in the LCA are typically set to zeros (Rozell et al., 2007), which can lead to slower convergence, especially under $\ell_{0}$ constraints due to the non-convexity of the optimization landscape. This issue underscores the importance of effective initialization strategies to accelerate convergence when employing $\ell_{0}$ regularization in LCA.
@@
Recent implementations of the LCA on neuromorphic hardware have alleviated this problem to a certain extent by significantly enhancing the computational efficiency and execution speed by several orders of magnitude (Davies et al., 2021; Henke et al., 2022; Parpart et al., 2023). Despite this progress, the algorithm still requires hundreds of optimization iterations and encounters substantial challenges with local optima. Thus, even with neuromorphic hardware, there is significant room for improvement in the underlying algorithmic principles.
@@ 2.3 Warm Starting with Deep Learning
Sambharya et al. (2023) introduced a framework for learning warm starts specifically for Douglas-Rachford splitting to solve convex quadratic programs (QPs), providing generalization guarantees for a broad range of operators. Similarly, Klaučo et al. (2019) used neural networks to warm-start active set methods in model predictive control (MPC), demonstrating significant reductions in computation time.
Beyond convex optimization, deep learning has been applied to accelerate non-convex optimization algorithms. For example, Sjölund and Bånkestad (2022) used graph neural networks to expedite matrix factorization, and developed schemes for quickly solving fixed-point problems. Inverse problems, such as sparse coding ((Gregor and LeCun, 2010), (Xin et al., 2016)), image restoration (Rick Chang et al., 2017), and wireless communication (He et al., 2020), have also benefited from embedding algorithm steps into deep networks. A widely used technique is unrolling algorithmic steps, which differentiates through these steps to minimize a performance loss (Monga et al., 2021; Diamond et al., 2017; Gregor and LeCun, 2010).
@@ 3.1 Predictor Network and Integration with LCA
The primary function of the predictor network is to provide an initial guess for the states $u_{i}$ of the LCA, which are then refined through iterative optimization. This approach addresses the inefficiencies inherent in traditional LCA, which often requires numerous iterations to converge. Inspired by the findings of Xin et al. (2016), which demonstrate deep networks’ ability to recover minimal $\ell_{0}$-norm representations in high dimensional scenarios, our methodology strategically employs a predictive model to initialize the LCA.
@@
Note that for this work we did not further explore the full range of possible CNN architecture for the predictor network, but settled on a trade-off between model complexity and effectiveness (see Table 4 for more details on the architecture). What we found in preliminary experiments, however, is that a subtractive ReLU arrangement as used in Lang et al. (2023), combining the computation branches in the network was necessary to balance sparsity and richness of the output:
${u}_{i}=\sigma(\operatorname{relu({a}_{i})}-\operatorname{relu}({b}_{i}))$ (6)
where $a_{i}$ and $b_{i}$ are output of the computational paths of the architecture shown in Table 4 for a single kernel $i$ and $\sigma$ is the standard sigmoid function. This setup allows to establish a stable reference value that is neither the minimum nor the maximum and also the branches to specialize on the range above and below this reference value, respectively.

### v2

@@ Abstract
The locally competitive algorithm (LCA) can solve sparse coding problems across a wide range of use cases. Recently, convolution-based LCA approaches have been shown to be highly effective for enhancing robustness for image recognition tasks in vision pipelines. To additionally maximize representational sparsity, LCA with hard-thresholding can be applied. While this combination often yields very good solutions satisfying an $\ell_{0}$ sparsity criterion, it comes with significant drawbacks for practical application: (i) LCA is very inefficient, typically requiring hundreds of optimization cycles for convergence; (ii) the use of a hard-thresholding results in a non-convex loss function, which might lead to suboptimal minima. To address these issues, we propose the Locally Competitive Algorithm with State Warm-up via Predictive Priming (WARP-LCA), which leverages a predictor network to provide a suitable initial guess of the LCA state based on the current input. Our approach significantly improves both convergence speed and the quality of solutions, while maintaining and even enhancing the overall strengths of LCA. We demonstrate that WARP-LCA converges faster by orders of magnitude and reaches better minima compared to conventional LCA. Moreover, the learned representations are more sparse and exhibit superior properties in terms of reconstruction and denoising quality as well as robustness when applied in deep recognition pipelines. Furthermore, we apply WARP-LCA to image denoising tasks, showcasing its robustness and practical effectiveness. Our findings confirm that the naive use of LCA with hard-thresholding results in suboptimal minima, whereas initializing LCA with a predictive guess results in much better outcomes.
keywords:
Computer Vision , Convolutional Sparse Coding , Locally Competitive Algorithms
journal: Neurocomputing
[…]
[inst1]organization=Mercedes-Benz AG, addressline=Leibnizstraße 2, city=Böblingen, country=Germany
[…]
[inst2]organization=Institut für Robotik und Kognitive Systeme, Universität zu Lübeck, addressline=Ratzeburger Allee 160, city=Lübeck, country=Germany
{graphicalabstract}
{highlights}
WARP-LCA accelerates convergence and achieves superior sparsity compared to LCA.
Achieves higher PSNR and SSIM with fewer iterations than traditional LCA.
Improves denoising in classification pipelines under varying noise levels.
Enables generalizable and efficient sparse coding with predictive initialization.
@@ 1 Introduction
Sparse coding has deep roots in neuroscience. The idea that sensory systems learn and represent the statistics of natural scenes date back to Barlow’s efficient coding hypothesis [4]. Barlow reasoned that redundancy reduction is key, since information retrieval in the sensory stream is akin to finding the needle in the haystack. Sparse coding then originated from the insight that sparse, redundant representations may be useful to make the statistical components of sensory information explicit [16, 3]. Within this context, the sparse coding algorithm, introduced by [38], has emerged as a paradigm to successfully model and predict response patterns of the primary visual cortex [27, 28, 6, 15].
@@
The human cortex is a very homogeneous structure, and it has been hypothesized that it performs the same operation everywhere [2]. The claim that sparse coding, at least to some degree, models general cortical function, reflects in a manifold of applications in which the algorithm has proven useful. It has early on been shown that the learned features are useful for inference [43], especially when unsupervised learning can be leveraged, like in stereo vision [33, 32, 15] or in scenarios with poorly labeled data [60]. The capabilities for unsupervised learning are most prominent in multi layer sparse coding networks, which are capable of learning abstract concepts [22, 59, 13]. In addition, sparse coding has proven at least state-of-the-art performance in removing noise and other corruptions from images [26]. More recently, Teti et al. [50] have shown that applying sparse coding as a pre-processing stage in image recognition pipelines yields state-of-the-art robustness to common corruptions and adversarial attacks.
@@
A significant challenge for applications with sparse coding is the computational expense of the inference algorithm. Finding sparse coefficients for each input sample requires an expensive iterative optimization process via gradient descent. One approach to resolve poor execution speed is the implementation on neuromorphic hardware. Rozell et al. [45] introduced the Locally Competitive Algorithm (LCA) as a biologically plausible model for sparse coding. Indeed, the implementation of the LCA on Intel Loihi has proven the strongest gain in energy efficiency and execution speed against the implementation on conventional hardware, as compared to any other implemented algorithm [10, 19, 41]. Considerable research has been directed towards developing efficient optimization algorithms for sparse coding on conventional hardware. Notable contributions include the works of Li and Osher [30], Mairal et al. [34], Beck and Teboulle [5], and Vonesch and Unser [52]. Gregor and LeCun [17] contributed significantly to the field by introducing the Learned ISTA (LISTA) algorithm, which has spawned a family of model-based sparse encoders (notable models include AMP-Net Zhang et al. [61], FISTA-Net Xiang et al. [53] and ISTA-Net Zhang and Ghanem [58].
@@
Sparse coding is typically associated with $\ell_{1}$ norm regularization due to its convexity, which facilitates a tractable optimization process [62]. However, for achieving maximum sparsity, $\ell_{0}$-like regularization is preferred [39]. Despite its non-convex nature, which introduces computational challenges, $\ell_{0}$ regularization directly targets sparsity in the solution [36]. Nonetheless, the Locally Competitive Algorithm (LCA) presents a distinctive case. While the discrete version of LCA implementing $\ell_{1}$ sparsity aligns mathematically with ISTA  [1], employing an $\ell_{0}$-like cost function in LCA diverges significantly in outcomes when compared to greedy methods such as Basis Pursuit (BP) [44, 45].
@@
Our main contribution of this paper is the development and validation of the Locally Competitive Algorithm with State Warm-up via Predictive Priming (WARP-LCA). This novel approach leverages a predictor network to provide an initial guess of the LCA state based on the current input, significantly improving convergence speed and solution quality. WARP-LCA addresses the inefficiencies and suboptimal solutions associated with LCA using hard-thresholding ($\ell_{0}$). By incorporating predictive priming, WARP-LCA dramatically reduces the number of optimization cycles required for convergence. Furthermore, WARP-LCA achieves superior minima, yielding more sparse and robust representations that enhance reconstruction and denoising quality, particularly in deep recognition pipelines. Our experiments, in the domain of image denoising and comparing WARP-LCA directly to LCA, confirm that WARP-LCA not only overcomes the limitations of traditional hard-thresholding LCA but also advances the field of biologically inspired deep learning by providing a novel and efficient method for convolutional sparse coding.
@@ 2.1 Convolutional Sparse Coding
The inference process involves estimating the sparse coefficients $a$ for a given data point $x$. This is typically achieved by minimizing an energy function that combines a reconstruction error term with a sparsity-inducing regularization term. A common formulation uses $\ell_{p}$ regularization to enforce sparsity [51].
@@
Learning in convolutional sparse coding involves updating the convolutional filters $\Phi$ based on the inferred sparse coefficients across a batch of data points. This is typically done through gradient descent on the energy function with respect to $\Phi$. Although the joint optimization of $\Phi$ and $a$ is non-convex, iterative approaches often converge to useful solutions, especially in natural signal domains. In addition to convolutional sparse coding, other widely known approaches for learning dictionaries include Principal Component Analysis (PCA) and Independent Component Analysis (ICA). PCA learns an orthogonal dictionary that captures the directions of maximum variance in the data. ICA, on the other hand, seeks to represent data as linear combinations of statistically independent components, which is useful for blind source separation and feature extraction [21].
@@ 2.2 Locally Competitive Algorithm (LCA)
The Locally Competitive Algorithm (LCA), introduced by Rozell et al. [45], represents a biologically plausible approach to sparse coding, where the dynamics of the neuronal population can be conceptualized as a system of coupled differential equations that govern the temporal evolution of neuronal activities. Unlike iterative thresholding methods or matching pursuit, the LCA exhibits a ’charging circuit behavior’ where neurons dynamically compete to represent the input.
@@
Rozell et al. [45] introduced a generalized thresholding function that enables the Locally Competitive Algorithm (LCA) to tackle various sparse coding problems by adjusting its parameters. The generalized thresholding function is expressed as:
@@
Specifically, for $\ell_{0}$ regularization—which aims to minimize the number of non-zero coefficients—the thresholding function approaches a hard threshold as $\gamma$ becomes large. This effectively turns the generalized thresholding function into the well-known proximal operator:
@@
The initial values of the states in the LCA are typically set to zeros [44], which can lead to slower convergence, especially under $\ell_{0}$ constraints due to the non-convexity of the optimization landscape. This issue underscores the importance of effective initialization strategies to accelerate convergence when employing $\ell_{0}$ regularization in LCA.
@@
Recent implementations of the LCA on neuromorphic hardware have alleviated this problem to a certain extent by significantly enhancing the computational efficiency and execution speed by several orders of magnitude [10, 19, 41]. Despite this progress, the algorithm still requires hundreds of optimization iterations and encounters substantial challenges with local optima. Thus, even with neuromorphic hardware, there is significant room for improvement in the underlying algorithmic principles.
@@ 2.3 Warm Starting with Deep Learning
Sambharya et al. [46] introduced a framework for learning warm starts specifically for Douglas-Rachford splitting to solve convex quadratic programs (QPs), providing generalization guarantees for a broad range of operators. Similarly, Klaučo et al. [23] used neural networks to warm start active set methods in model predictive control (MPC), demonstrating significant reductions in computation time.
Beyond convex optimization, deep learning has been applied to accelerate non-convex optimization algorithms. For example, Sjölund and Bånkestad [47] used graph neural networks to expedite matrix factorization, and developed schemes for quickly solving fixed-point problems. Inverse problems, such as sparse coding ([17], [55]), image restoration [42], and wireless communication [18], have also benefited from embedding algorithm steps into deep networks. A widely used technique is unrolling algorithmic steps, which differentiates through these steps to minimize a performance loss [35, 11, 17].
@@ 3 WARP-LCA
[…]
Figure 2: Illustration of the WARP-LCA Method. The WARP-LCA method integrates a fully convolutional neural network (CNN) to predict LCA states. These predicted states serve as a warm start for the LCA module. After several LCA iterations, the refined sparse activations are optionally processed through a transpose convolution block to reconstruct the image. For a detailed description see Table 4.
@@ 3.1 Predictor Network and Integration with LCA
The primary function of the predictor network is to provide an initial guess for the states $u_{i}$ of the LCA, which are then refined through iterative optimization. This approach addresses the inefficiencies inherent in traditional LCA, which often requires numerous iterations to converge. Inspired by the findings of Xin et al. [55], which demonstrate deep networks’ ability to recover minimal $\ell_{0}$-norm representations in high dimensional scenarios, our methodology strategically employs a predictive model to initialize the LCA.
@@
3.2 Designing and Training Networks to Predict LCA States
Our goal was to identify an effective trade-off between model complexity and prediction quality without exhaustively surveying the full range of possible CNN configurations (see Table 5 for more details). We investigated three main architectures across five size configurations: a simple feedforward network, a ResNet-inspired network, and a fully convolutional variant with dual branches and a subtractive arrangement in the output layer (titled WARP-CNN). Our preliminary experiments indicated that the WARP-CNN with a subtractive ReLu arrangement as described in Lang et al. [25] balanced the joined requirements on sparsity and accurate output in WARP-CNNs:
${u}_{i}=\sigma\bigl{(}\operatorname{relu}({a}_{i})-\operatorname{relu}({b}_{i})\bigr{)},$ (6)
where $a_{i}$ and $b_{i}$ are outputs of the two computation branches in the network and $\sigma$ denotes the sigmoid function. This construction establishes a stable reference that is neither the minimum nor the maximum, allowing the branches to specialize on ranges above and below this reference value, respectively.
Training employed the ADAM optimizer (learning rate $10^{-4}$). Smaller batch sizes consistently yielded lower losses, indicating that finer gradient updates helped avoid local minima. This arrangement reduced the loss landscape complexity and stabilized convergence. Moreover, the sparse codes used as targets were scaled to fall within $[0,1]$, to be compatible with the sigmoidal subtractive ReLU Activation as described above. Notably, the largest WARP-CNN model demonstrated superior performance among the tested setups (see 6 and 2 for a full description of the model architecture). Overfitting was largely absent for larger models, whereas smaller variants exhibited plateauing validation losses, suggesting limited representational capacity rather than typical overfitting. Such plateaus are consistent with the networks converging on convolutional kernel representations for image patches, as expected when capturing the underlying sparse code structures.
We introduce a custom loss function derived from a Laplacian prior, which can be interpreted as an approximate negative log-likelihood under a scale parameter $\beta_{i}=1+\epsilon+\gamma|t_{i}|$. Specifically, the Laplace log-density for a target $t_{i}$ given an output $o_{i}$ is
$-\log p(t_{i}\mid o_{i})\;\approx\;\log\bigl{(}\beta_{i}\bigr{)}+\frac{|t_{i}-o_{i}|}{\beta_{i}}.$
Although a Laplace prior typically implies an absolute deviation term, we adopt a squared-error surrogate for smoother gradients and more stable optimization, since the absolute-value function is not differentiable at zero. Since $\log\bigl{(}\beta_{i}\bigr{)}$ does not affect gradients with respect to model parameters it can be safely dropped for training objectives. Replacing the absolute deviation with a squared term yields:
$\mathcal{L}=\frac{1}{N}\sum_{i=1}^{N}\left((o_{i}-t_{i})^{2}\cdot\frac{1}{1+\epsilon+\gamma|t_{i}|}\right).$ (7)
Here:
  • 1.
    $N$ is the total number of elements,
  • 2.
    $o_{i}$ and $t_{i}$ are the model output and target,
  • 3.
    $\gamma$ is the Laplace scale parameter controlling how strongly to penalize deviations at small $t_{i}$,
  • 4.
    $\epsilon$ is a small constant for numerical stability.
In our experiments, this custom loss effectively prioritizes small, frequent values—critical for modeling sparse targets. A value of $\gamma=3$ proved consistently effective. Consequently, errors at small targets are multiplied by a larger factor (due to the smaller denominator), while errors at larger targets are down-weighted. Since the target values fall within $[0,1]$, $\gamma=3$ places roughly four times greater emphasis (ignoring $\epsilon$) on small $t_{i}$-values than on $t_{i}$ near one.